# Combinatorial flats and Schubert varieties of subspace arrangements

### v2

@@ 5. Problems
The Chow class of a Schubert variety of subspace arrangement $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ may not be multiplicity-free, and the coefficients $c_{\mathbf{b}}$ of $[Y_{V}]$ may depend on $\mathbf{n}$. We conjecture the following formula:
@@
If $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$, then when $\mathbf{b}$ is a basis, the coefficient of $[\mathbb{P}^{\mathbf{b}}]$ in $[Y_{V}]$ is $c_{\mathbf{b}}=\binom{n_{1}}{b_{1}}\binom{n_{2}}{b_{2}}\cdots\binom{n_{N}}{b_{N}}$.
5.5.
@@
An analogous question can be asked for the Kazhdan-Lusztig polynomial of [EPW16], which in the case of a realizable matroid corresponds to the local intersection cohomology of $Y_{V}$ at the point (,,)($\infty$,$\ldots$,$\infty$)( ∞ , … , ∞ ).

### v3

@@ 5. Problems
The Chow class of a Schubert variety of subspace arrangement $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ may not be multiplicity-free, and the coefficients $c_{\mathbf{b}}$ of $Y_{V}]$ may depend on $\mathbf{n}$. We conjecture the following formula:
@@
If $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$, then when $\mathbf{b}$ is a basis, the coefficient of $\mathbb{P}^{\mathbf{b}}]$ in $Y_{V}]$ is $c_{\mathbf{b}}=\binom{n_{1}}{b_{1}}\binom{n_{2}}{b_{2}}\cdots\binom{n_{N}}{b_{N}}$.
$5.$
@@
An analogous question can be asked for the Kazhdan-Lusztig polynomial of [EPW16], which in the case of a realizable matroid corresponds to the local intersection cohomology of $Y_{V}$ at the point $(\infty,\ldots,\infty)$.